Three charges of magnitude 3.0 x10-4 C each are located atx = 1.0 m, y = 0.0 m,at x = 0.0 m, y =0.0 m and at x =-1.0 m, y = 0.0 m. The one in themiddle is negative, while the other two are positive. What is thenet Coulombic force exerted by them on a negative 3.0x 10-5 C charge located atx = 0.0 m, y = 2.0 m? (The value of k is 9.0 x109 N*m2/C2.)

a.
87 N attractive
b.
87 N repulsive
c.
120 N repulsive
d.
120 N attractive
Answer #1
The hypoteneus of the right Δ,
distance of the charge above from the end charges)is =√(1 +4) = √5 cm
The forces,
         F1 = F2 =9x109x3x3x10-9/5
                     = 16.2 N attraction
The force F3 =9x109x3x3x10-9/4
                     = 20.25 N repulsion
Resolve into rectangular components along y axis and xaxis
The x components cancel out.
The resultant force, R =2.F1 Cos θ - F3,
where Cos θ =2/√5 = 0.8944.
Hence,
                R = 2x16.2x0.8944 - 20.25 N
                     =28.98 - 20.25
                    = 8.73 N toward the origin.
This is attractive force.
But the Choice (a) is 87 N attractive.
I think my answer is correct.